# Long-Context Generalization with Sparse Attention

### v2

@@ Discussion.
The results, shown in Table 5, reveal two critical factors for length generalization: positional encoding and attention sparsity. With RoPE, all methods struggle beyond 2$\times$ training length, while NAPE (NoPE + ALiBi) enables substantially better generalization across all tasks. Within NAPE variants, ASEntmax dramatically outperforms others at extreme lengths—maintaining 96.4% accuracy at 256$\times$ test length on MQMTAR (vs. 80.2% for softmax) and 96.4% at 4$\times$ on Reverse (vs. 0% for softmax). This positional encoding advantage aligns with our theoretical analysis showing that $\alpha$-entmax transforms positional biases into adaptive attention windows.
Task-specific patterns further validate our theoretical claims. The consistent superiority of ASEntmax over basic $\alpha$-entmax confirms the benefits of adaptive scaling, particularly at extreme lengths where fixed-$\alpha$ may become too sparse or too diffuse. SSMax performs well on the Copy task, even outperforming other methods, but struggles on more complex tasks like MQMTAR and Reverse at extreme lengths. This indicates that while scaling logits helps maintain peak attention magnitude, the explicit sparsity of $\alpha$-entmax provides additional benefits by completely removing irrelevant connections. These findings are further supported by results on the Max Retrieval task (Figure 5), where sparse attention mechanisms demonstrate superior length extrapolation compared to dense approaches, with ASEntmax maintaining over 60% accuracy even at 4096-length sequences—a dramatic improvement over standard Softmax and Adaptive Temperature (Veličković et al., 2024).
Finally, Copy and Reverse show moderate generalization (up to 64$\times$ and 8$\times$ respectively), while Sort fails beyond 4$\times$ length for all methods. This pattern suggests that tasks requiring precise global ordering (Sort, Reverse) are inherently more challenging for length generalization than tasks dependent on local or independent token properties. We provide per-task results, including results for other positional encoding methods such as ALiBi and NoPE, in §H.
@@
Method ID 2×2$\times$2 × 4×4$\times$4 × 16×16$\times$16 × 64×64$\times$64 × 256×256$\times$256 × 1024×1024$\times$1024 × ID 1.5×1.5$\times$1.5 × 2×2$\times$2 × 4×4$\times$4 × 8×8$\times$8 ×
RoPE Softmax 100.0 3.1 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0
SSMax 99.8 6.2 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0
Entmax 99.8 49.4 4.5 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0
ASEntmax 100.0 66.9 0.8 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0
[…]
SSMax 99.9 100.0 99.9 99.4 94.1 53.4 1.0 100.0 54.6 0.0 0.0 0.0
@@
ID 2×2$\times$2 × 4×4$\times$4 × 8×8$\times$8 × 16×16$\times$16 × 32×32$\times$32 × 64×64$\times$64 × ID 2×2$\times$2 × 4×4$\times$4 × 8×8$\times$8 × -
RoPE Softmax 100.0 2.8 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 -
SSMax 100.0 0.0 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 -
Entmax 100.0 34.3 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 -
ASEntmax 100.0 5.3 0.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 -
[…]
@@
Figure 5: Accuracy (%) on the Max Retrieval task across different sequence lengths. Adaptive Temperature (Adapt. Temp.) represents the approach proposed by Veličković et al. (2024).
@@ 6 Related Works
Understanding and improving length generalization is a very active research area. We summarize below the related work in different dimensions.
@@
Recent work has identified attention dispersion as a fundamental limitation in softmax-based transformers (Dong et al., 2021; Zhai et al., 2023; Veličković et al., 2024). For example, Veličković et al. (2024) demonstrate that softmax attention inevitably disperses focus as sequence length increases, while Nakanishi (2025) propose SSMax to scale attention logits based on sequence length. Our approach employs $\alpha$-entmax (Peters et al., 2019), which naturally produces sparse distributions by assigning exactly zero probability to irrelevant tokens. We provide theoretical guarantees that $\alpha$-entmax maintains bounded normalized entropy as sequence length increases—a property softmax fundamentally lacks. Our ASEntmax further improves $\alpha$-entmax with learnable, context-dependent scaling, leading to consistent gains over SSMax across diverse tasks.
@@
The interaction between attention mechanisms and positional encodings significantly impacts long-context generalization. Wu et al. (2025) show that causal masking with distance bias creates a position advantage for early tokens, while Barbero et al. (2025) analyze frequency components in RoPE (Su et al., 2024). Our work shows how $\alpha$-entmax transforms these biases: with NoPE, it helps attention to not always converge to early tokens; with ALiBi, it creates hard attention windows—similar to Hard-ALiBi (Jelassi et al., 2024); with RoPE, it induces frequency-dependent sparsity. Based on this analysis, we introduce NAPE (NoPE + ALiBi), which combines ALiBi’s consistent recency bias with NoPE’s flexible content-based attention for improved long-context generalization. Similar to NAPE, Fu et al. (2025) propose to combine heads with ALiBi and RoPE to increase training stability.
@@
Previous sparse attention approaches include structured patterns like Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), Flash-Attention (Dao et al., 2022), as well as adaptive methods like $\alpha$-entmax (Peters et al., 2019) and Sparsefinder (Treviso et al., 2022), recently accelerated by AdaSplash (Gonçalves et al., 2025). Our work differs by providing a theoretical analysis of why sparsity helps with long-context generalization, linking it to fundamental limitations including dispersion, representational collapse, and over-squashing.
@@ 7 Conclusions
In this paper, we present a principled approach to long-context modeling by replacing softmax with $\alpha$-entmax in transformer attention. Our theoretical analysis demonstrates how this simple change addresses three fundamental limitations: it avoids attention dispersion through naturally sparse distributions, prevents representational collapse by maintaining distinct token representations, and alleviates over-squashing by reducing gradient paths from $\mathcal{O}(n^{L})$ to $\mathcal{O}(s^{L})$, where $s\ll n$ is the number of tokens with nonzero probability. We further introduce Adaptive-Scalable $\alpha$-entmax (ASEntmax), which adaptively adjusts sparsity based on sequence length for each attention head and query input. Our empirical results confirm these theoretical predictions, with ASEntmax significantly outperforming softmax and existing alternatives across diverse tasks, maintaining high performance even at sequence lengths 1000$\times$ beyond training examples. These findings suggest that addressing the fundamental mathematical limitations of transformer attention mechanisms may provide a direct path to achieve robust long-context generalization.
Limitations.
While our experiments focus on controlled settings with small transformers, the theoretical principles we establish should extend to deeper architectures and larger scales; however, future research on sparse attention in production-scale models is still needed, especially with multi-stage approaches commonly employed for extending the context-length in traditional LLMs (Grattafiori et al., 2024).
@@ Acknowledgments
We thank Hugo Pitorro for his insightful and constructive comments throughout this work. We also thank the SARDINE Lab members for reviewing this paper and providing helpful feedback. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for ResponsibleAI), by the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
@@ Appendix A $\alpha$-entmax Transformation
where $[\cdot]_{+}=\max(0,\cdot)$, and $\tau(\bm{z})$ is chosen so that $\bm{p}^{\star}$ sums to $1$. Figure 2 illustrates how $\alpha$-entmax with tempered scores $\bm{z}/\theta$ behaves for different choices of $\alpha$.
@@ Appendix B Model Definition and Notation
In our approach, we replace softmax with $\alpha$-entmax (§A). We group the attention weights into an attention matrix at the $\ell$-th layer, defined element-wise as $[\bm{P}^{(\ell)}]_{ij}:=p_{ij}^{(\ell)}$. This is a row-stochastic lower triangular matrix that can also be interpreted as a probabilistic directed graph. Finally, when incorporating positional information, we modify the attention logits computation according to the chosen positional encoding strategy:
@@
Appendix C Representational Collapse and Over-squashing
C.1 Proof of Lemma 3.1
For $\alpha>1$, we need to find the value $b_{\max}$ such that, for any $b\leq b_{\max}$, $\operatorname{\alpha\text{-}entmax}(\bm{x}^{*})_{n}=0$ holds. From the definition of $\alpha$-entmax (2), a token $i$ receives non-zero probability iff $(\alpha-1)z_{i}>\tau(\bm{z})$, where $\tau(\bm{z})$ is the threshold ensuring the sum of probabilities equals 1. Therefore, for the token $b$ in the extended sequence $\bm{x}^{*}$ to receive zero probability (thus not affecting other probabilities), we need:
$(\alpha-1)b\leq\tau(\bm{x}^{*}).$ (18)
$(\alpha-1)b\leq\tau(\bm{x}).$ (19)
Solving for $b$, we get:
$b\leq\frac{\tau(\bm{x})}{\alpha-1}.$ (20)
Thus, we can define $b_{\max}=\frac{\tau(\bm{x})}{\alpha-1}$. For any $b\leq b_{\max}$, the token at position $n$ in $\bm{x}^{*}$ receives zero attention, meaning it doesn’t affect the normalization. Therefore, $\tau(\bm{x}^{*})=\tau(\bm{x})$, which means that the condition (20) is both necessary and sufficient, and:
C.2 Proof of Proposition C.2
We prove Proposition C.2 by exhibiting the counterexample below.
{proposition}
where the final token $\bm{v}_{a}\in\mathbb{R}^{d}$ is repeated.
$\|\bm{v}_{n}^{(L)}-\bm{v}_{n+1}^{*(L)}\|_{1}\geq c>0$ (23)
Proof.
We prove this through explicit construction.
Consider the specific embedding choice where:
This construction yields the following attention logits:
$\bm{z}_{n}^{(1)}$ $=[b,b,\ldots,b,\phi],$ (26)
$\bm{z}_{n+1}^{*(1)}$ $=[b,b,\ldots,b,\phi,\phi].$ (27)
Specific counterexample.
$\bm{z}_{n}^{(1)}$ $=[0,\ldots,0,0.5],$ (28)
$\bm{z}_{n+1}^{*(1)}$ $=[0,0,\ldots,0,0.5,0.5].$ (29)
For $\alpha=2.0$ (sparsemax), the attention distributions are:
[…]
where $10<p_{b}<p_{n}<$ 1. As $n\to\infty$, we have $p_{n}\to p^{*}$ where:
This demonstrates that the $L_{1}$ difference remains bounded at approximately $c=0.354$, independent of sequence length $n$. This specific example already establishes the existence of a counterexample to representational collapse with $\alpha$-entmax. We now extend this result to prove Proposition C.2 for general constructions of $\bm{v}$ and $\bm{v}_{a}$.
General Construction.
With values of $\phi$ such that $b+2^{-(\alpha-1)}/(\alpha-1)\leq\phi<b+1/(\alpha-1)$, $\alpha$-entmax produces the following distributions:333 The upper bound ensures a dense output for $\bm{z}_{n}^{(1)}$, following Lemma 3.1 with $\tau=(\alpha-1)\phi-1$. The lower bounds ensure a sparse output for $\bm{z}_{n+1}^{*(1)}$, following Lemma D with $k=2$.
In particular, since the first $n-1$ positions share the same representation, we have $p^{(1)}_{1}=p^{(1)}_{2}=...=p^{(1)}_{n-1}=(1-p^{(1)}_{n})/(n-1)=p^{(1)}_{b}>0$, with $10<p^{(1)}_{n}<$ 1. This leads to representations:
$\bm{v}_{n+1}^{*(1)}$ $=\frac{1}{2}\bm{v}_{a}+\frac{1}{2}\bm{v}_{a}=\bm{v}_{a}.$ (37)
$=(1-p_{n}^{(1)})\left\|\bar{\bm{v}}-\bm{v}_{a}\right\|_{1}.$ (39)
First condition.
We need to choose parameters so that as $n\to\infty$, the original sequences remains dense and the extended sequence is in the sparse regime. From our analysis with Lemma 3.1 and D, this requires:
$\frac{2^{-(\alpha-1)}}{\alpha-1}\leq\phi-b<\frac{1}{\alpha-1}.$ (40)
From the $\alpha$-entmax definition, we have:
$(p_{b}^{(1)})^{\alpha-1}$ $=(\alpha-1)b-\tau,$ (41)
$(p_{n}^{(1)})^{\alpha-1}$ $=(\alpha-1)\phi-\tau,$ (42)
From the normalization constraint $(n-1)p_{b}^{(1)}+p_{n}^{(1)}=1$:
$p_{b}^{(1)}=\frac{1-p_{n}^{(1)}}{n-1}.$ (44)
Substituting:
We know from the dense regime that $p_{n}^{(1)}\to p^{*}$, with $10<p^{*}<$ 1. Thus, as $n\to\infty$, the probability $p_{n}^{(1)}$ satisfies:
$(p^{*})^{\alpha-1}=(\alpha-1)(\phi-b)$ (47)
Therefore:
$p^{*}=[(\alpha-1)(\phi-b)]^{\frac{1}{\alpha-1}}<1.$ (48)
Second condition.
  • $\bm{v}_{i}=\frac{b}{\sqrt{\phi}}\bm{e}_{1}+\bm{e}_{2}$ for $i=1,\ldots,n-1$.
Note that this construction satisfies the logit constraints:
$\bm{v}_{a}^{T}\bm{v}_{i}$ $=\sqrt{\phi}\cdot\frac{b}{\sqrt{\phi}}+0\cdot 1=b,$ (49)
$\bm{v}_{a}^{T}\bm{v}_{a}$ $=(\sqrt{\phi})^{2}=\phi.$ (50)
The average representation is:
Since $\bm{v}_{a}=\sqrt{\phi}\bm{e}_{1}$, we have:
The bound is strictly positive because $b\neq\phi$ by the logit difference requirement, and because of the constant $+1$ term from the $\bm{e}_{2}$ component. Therefore, $\lim_{n\to\infty}\|\bar{\bm{v}}-\bm{v}_{a}\|_{1}\not\to 0$. For the case $d=1$, we can use the simpler construction $v_{i}=\frac{b}{\sqrt{\phi}}$ and $v_{a}=\sqrt{\phi}$, where the non-collapse condition becomes verifying that $\frac{b}{\sqrt{\phi}}\neq\sqrt{\phi}$, which follows from $b\neq\phi$. In contrast, as shown by Barbero et al. (2024), the resulting representations become increasingly similar as $n\to\infty$ with softmax ($\alpha=1.0$), regardless of the input content, leading to representational collapse. ∎
[Uncaptioned image][Uncaptioned image]
Empirical Verification of Representational Preservation.
  1. 1.
    Constant prefixes: $b=1$ and $\phi=1.2$.
  2. 2.
    Random prefixes: $b\sim\mathcal{U}(0,1)$ and $\phi=1.2$.
C.3 Proof of Proposition 3.3
{proposition}
$\bm{y}_{n}=\mathrm{norm}_{3}\bigl{(}\bm{v}_{n}^{(L)}\bigr{)}$ (53)
Proof.
$\mathcal{T}_{1}$ =$=\{k\in\{i,i+1,\ldots,n\}:k\in\mathcal{S}_{i}^{(0)}\text{ or }k=i\},$ (60)
$\mathcal{T}_{\ell}(k_{\ell-1})$ =$=\{k\in\{k_{\ell-1},k_{\ell-1}+1,\ldots,n\}:k\in\mathcal{S}_{k_{\ell-1}}^{(\ell-1)}\text{ or }k=k_{\ell-1}\}\quad\text{for }\ell>1.$ (61)
These sets have the following meaning:
Empirical Verification of Over-squashing Alleviation.
Appendix D Non-Dispersion of $\alpha$-entmax

### v3

@@ Discussion.
[…]
@@
Method ID $2\times$ $4\times$ $16\times$ $64\times$ $256\times$ $1024\times$ ID $1.5\times$ $2\times$ $4\times$ $8\times$
Softmax 100.0 100.0 100.0 99.5 97.8 80.2 3.0 100.0 36.0 0.0 0.0 0.0
Top-K (K=32) 100.0 99.9 6.2 0.0 0.0 0.0 0.0 100.0 100.0 98.7 57.0 0.0
SSMax 99.9 100.0 99.9 99.6 98.3 90.6 26.7 100.0 54.6 0.0 0.0 0.0
@@
ID $2\times$ $4\times$ $8\times$ $16\times$ $32\times$ $64\times$ ID $2\times$ $4\times$ $8\times$ -
Softmax 100.0 100.0 99.9 99.9 99.4 96.1 85.5 100.0 0.0 0.0 0.0 -
Top-K (K=32) 100.0 99.7 96.8 26.7 0.0 0.0 0.0 100.0 92.5 0.0 0.0 -
@@
5.2 Language Modeling
To validate our approach on real-world tasks, we train 420420M-parameter decoder-only models following the LLaMA 3 architecture (details in App. G.3) on the high-quality DCLM-Edu dataset (Allal et al., 2025), for 77B tokens with a context length of $n=2048$. As in the synthetic experiments, we report results in the main paper using NAPE (NoPE+ALiBi). Results with RoPE are reported in App. H.9. We evaluate short-context performance on Lambada, HellaSwag, PIQA, ARC-C, Winogrande, and OpenBookQA, and assess long-context generalization via perplexity on ArXiv/PubMed subsets of The Pile (Gao et al., 2020) as well as RULER’s Needle-in-a-Haystack tasks (Hsieh et al., 2024).
Table 2: Downstream-task results on short-context datasets. Best results are in bold.
Method Lambada (PPL) Lambada Hellaswag PIQA Arc-C WinoGrande OpenbookQA
Softmax 52.4 30.9 33.1 65.1 25.6 49.5 28.2
SSMax 48.9 31.6 32.9 65.1 25.0 51.5 30.4
Entmax 47.9 32.1 32.8 63.6 24.6 50.9 29.0
ASEntmax 41.6 34.3 33.4 63.8 26.0 50.0 28.6
Table 3: Long-context perplexity on ArXiv and PubMed from The Pile. Best results are in bold.
ArXiv (ID) ArXiv (OOD) PubMed (ID) PubMed (OOD)
Model 1K 2K 4K 8K 16K 1K 2K 4K 8K 16K
Softmax 20.82 16.47 13.74 12.69 12.70 18.36 17.34 15.31 15.61 18.73
SSMax 20.81 16.44 13.41 11.95 12.00 18.13 17.10 14.87 14.11 15.36
Entmax 23.40 18.57 15.36 14.02 14.68 22.33 21.14 18.94 18.31 20.29
ASEntmax 20.64 16.30 13.15 11.19 10.77 18.15 17.09 14.66 13.48 14.48
Table 4: Retrieval performance on RULER benchmark. All models use NAPE positional encoding and were trained on 2048-token contexts. Best results are in bold.
S-NIAH-1 (ID) S-NIAH-1 (OOD) S-NIAH-2 (ID) S-NIAH-2 (OOD)
Model 1K 2K 4K 8K 16K 1K 2K 4K 8K
Softmax 100.0 99.4 94.2 11.4 0.8 100.0 100.0 4.8 0.0
SSMax 100.0 99.8 99.2 92.0 75.2 99.4 99.2 64.4 14.8
Entmax 99.8 99.8 89.0 21.6 1.2 99.6 99.4 64.8 7.2
ASEntmax 99.6 100.0 100.0 99.8 97.4 99.4 99.4 83.2 25.4
Discussion.
In short-context evaluation (Table 2), all models perform comparably. We note that performance on ARC-C, Winogrande, and OpenBookQA is near-random. Among the other benchmarks, ASEntmax achieves the highest scores on Lambada (both in terms of perplexity and accuracy) and HellaSwag, while Softmax and SSMax perform best on PIQA. For long-context modeling (Table 3), ASEntmax outperforms all other models. On ArXiv, it shows strong extrapolation, maintaining a decreasing perplexity trend even when extended to $8\times$ the pre-trained sequence length. On PubMed, although all models struggle with $8\times$ extrapolation, ASEntmax still leads by a margin of about 1 perplexity point vs SSMax. In retrieval tasks from RULER (Table 4), entmax-based models perform better overall, with both Entmax and ASEntmax surpassing their softmax counterparts at the extrapolation $4\times$ and $8\times$. ASEntmax in particular shows strong extrapolation in passkey retrieval on the simple haystack task (S-NIAH-1), maintaining near-perfect performance even at $8\times$ context length. On the harder S-NIAH-2 variant, despite performance dropping substantially across all models, ASEntmax remains the best performer at $2\times$ and $4\times$ length generalization.
@@ 6 Related Works
Recent work has identified attention dispersion as a fundamental limitation in softmax-based transformers (Dong et al., 2021; Zhai et al., 2023; Veličković et al., 2025). For example, Veličković et al. (2025) demonstrate that softmax attention inevitably disperses focus as sequence length increases, while Nakanishi (2025) propose SSMax to scale attention logits based on sequence length. Our approach employs $\alpha$-entmax (Peters et al., 2019), which naturally produces sparse distributions by assigning exactly zero probability to irrelevant tokens. We provide theoretical guarantees that $\alpha$-entmax maintains bounded normalized entropy as sequence length increases—a property softmax fundamentally lacks. Our ASEntmax further improves $\alpha$-entmax with learnable, context-dependent scaling, leading to consistent gains over SSMax across diverse tasks.
@@
The design of positional encodings plays a central role in enabling transformers to generalize to long contexts. ALiBi (Press et al., 2022) introduced linear attention biases with fixed slopes that encourage recency, and has since inspired several extensions that either mitigate softmax-related issues (e.g., Hard-ALiBi (Jelassi et al., 2024)) or induce various learned decay patterns (KERPLE (Chi et al., 2022), FIRE (Li et al., 2024)). More recently, Stick-Breaking Attention (Tan et al., 2025) can be interpreted as a dynamic variant of ALiBi, in which the slopes are input-dependent and calculated adaptively from spans of tokens. Our work incorporates NAPE (NoPE + ALiBi). In this formulation, ALiBi heads combined with $\alpha$-entmax create hard attention windows similar to Hard-ALiBi (Jelassi et al., 2024). NoPE, in turn, can learn positional bias (see App. H.1), and when combined with ASEntmax, it induces a learnable, input-dependent recency bias.
@@
Previous sparse attention approaches include structured patterns like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), as well as adaptive methods like $\alpha$-entmax (Peters et al., 2019; Gonçalves et al., 2025), top-$k$ related methods (Gupta et al., 2021; Treviso et al., 2022; Zeng et al., 2025), and chunk-based approaches (Mohtashami & Jaggi, 2023; Hu et al., 2025). Our work differs by providing a theoretical analysis of why sparsity helps with long-context generalization, linking it to fundamental limitations including dispersion, representational collapse, and over-squashing. In particular, we note chunk-based approaches are orthogonal to our method and can be integrated with them to achieve even stronger length generalization and possibly faster inference.
@@ 7 Conclusions
[…]
Reproducibility Statement
To ensure reproducibility of our results, our complete implementation and experimental configurations will be publicly available at https://github.com/deep-spin/asentmax. For efficient softmax attention, we have used FlashAttention-2 (Dao, 2024), and for $\alpha$-entmax we relied on AdaSplash (Gonçalves et al., 2025). All theoretical results presented in §3 are accompanied by complete proofs in Appendix §C and §D. For our synthetic task experiments, we provide detailed task descriptions, data generation procedures, model architectures, and hyperparameters in Appendix §G.1-G.2. Our language modeling experiments use publicly available datasets: DCLM-Edu (Allal et al., 2025) for training and standard benchmarks (LAMBADA, HellaSwag, PIQA, ARC-C, Winogrande, OpenBookQA), The Pile (Gao et al., 2020) (ArXiv, PubMed), and RULER (Hsieh et al., 2024) for evaluation, with complete model specifications and training details provided in Appendix §G.3.
@@ Acknowledgments
We thank the SARDINE Lab members for reviewing this paper and providing helpful feedback. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for ResponsibleAI), by the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
@@ Appendix A $\alpha$-entmax Transformation
where $\cdot]_{+}=\max(0,\cdot)$, and $\tau(\bm{z})$ is chosen so that $\bm{p}^{\star}$ sums to $1$. Figure 2 illustrates how $\alpha$-entmax with tempered scores $\bm{z}/\theta$ behaves for different choices of $\alpha$.
@@ Appendix B Model Definition and Notation
In our approach, we replace softmax with $\alpha$-entmax (§A). We group the attention weights into an attention matrix at the $\ell$-th layer, defined element-wise as $\bm{P}^{(\ell)}]_{ij}:=p_{ij}^{(\ell)}$. This is a row-stochastic lower triangular matrix that can also be interpreted as a probabilistic directed graph. Finally, when incorporating positional information, we modify the attention logits computation according to the chosen positional encoding strategy:
@@
Appendix C Non-Dispersion of $\alpha$-entmax